# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v1

@@ 3 Problem setting and preliminaries
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}&\displaystyle\mathbb{E}_{{\mathbb{P}}_{N}}\Bigg{[}\underset{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}{\sup}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}\Bigg{]}.\end{array}$ (8)
@@
8 reduces to 4 when $\alpha=0$. Note that 8 is identical to feature robust training [9] which is not motivated by adversarial attacks, but the presence of noisy observations in the training set [4, 29].
4 Wasserstein adversarially robust optimization
8 replaces the loss function of 4 with the worst-case loss (with respect to adversarial attacks). Here we show that 8 is equivalent to an ERM of a modified loss, which is convex and Lipschitz.
Proposition 4.1.
Let $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y):=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ denote the adversarial loss associated with the logloss, and $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ its univariate counterpart. We have $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)=\sup_{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}$ and so 8 is identical to
@@ 4 Wasserstein adversarially robust optimization
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}&\displaystyle\mathbb{E}_{{\mathbb{P}}_{N}}[\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)].\end{array}$
Moreover, $\mathrm{Lip}(L^{\alpha})=1$ for any $\alpha\geq 0$.
The proof is in Appendix B.1. Proposition 4.1 tells us that true expected loss under adversarial attacks is $\mathbb{E}_{\mathbb{P}^{0}}[\ell_{\bm{\beta}}^{\alpha}(\bm{x},y)]$. Therefore, instead of optimizing the empirical risk $\mathbb{E}_{\mathbb{P}_{N}}[\ell_{\bm{\beta}}(\bm{x},y)]$8 optimizes the empirical adversarial risk $\mathbb{E}_{\mathbb{P}_{N}}[\ell_{\bm{\beta}}^{\alpha}(\bm{x},y)]$. This means that 8 calibrates the loss function so that we train and test with the same loss $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)$. However, 8 still optimizes this loss for the empirical distribution ${\mathbb{P}_{N}}$ and is thus prone to overfitting due to the statistical error of estimating $\mathbb{P}^{0}$ with $\mathbb{P}_{N}$. To address overfitting in the adversarial setting (robust overfitting of 8), we derive a Wasserstein DR counterpart of 8. We start with the following assumption.
Assumption 4.2.
We are given finite $\varepsilon>0$ satisfying $\mathrm{W}(\mathbb{P}^{0},\mathbb{P}_{N})\leq\varepsilon$ (i.e., $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$).
We discuss relaxing this assumption in Section 6. We now introduce the distributionally and adversarially robust logistic regression problem:
@@
The following result shows that, for a fixed $\varepsilon$DR-ARO can be reformulated as a convex optimization problem. This is a direct corollary of Proposition 4.1 and Theorem 14 (ii) of [59]; see Appendix B.2.
Corollary 4.3.
@@
$\begin{array}[]{cll}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\bm{s}\end{subarray}}{\inf}&\varepsilon\lambda+\frac{1}{N}\sum_{i=1}^{N}s_{i}\\ \mathrm{s.t.}&\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},y^{i})\leq s_{i},\text{ }\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},-y^{i})-\lambda\kappa\leq s_{i}&\forall i\in[N]\\ &\lVert\bm{\beta}\rVert_{q^{\star}}\leq\lambda,\ \bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+}.\end{array}$
The constraints of this problem are exponential cone representable (Appendix C), and for $q\in\{1,2,\infty\}$, the yielding problem can be solved with the exponential cone solver of MOSEK [41] in polynomial time (with respect to their input size [44]). DR-ARO addresses the overfitting issue of 8 by solving its distributionally robust counterpart. However, the DRO approach of considering the worst-case distribution within a ball around the empirical distribution can be overly conservative [53]. Next, we explore how employing auxiliary data can reduce this conservatism.
5 Reducing conservatism of DR-ARO via intersection of Wasserstein balls
So far we have discussed the setting where we have access to an empirical distribution $\mathbb{P}_{N}$ that is constructed from $N$ i.i.d. samples of the true distribution $\mathbb{P}^{0}$. Suppose that we have an auxiliary distribution ^$\widehat{\mathbb{P}}_{\widehat{N}}$ which is constructed from $\widehat{N}$ i.i.d. samples $\{\widehat{\bm{\xi}}^{j}=(\widehat{\bm{x}}^{j},\widehat{y}^{j})\}_{j\in[\widehat{N}]}$ of another distribution $\widehat{\mathbb{P}}$. In this section, we explore how auxiliary data can help us identify a subset of the Wasserstein ball $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ in which $\mathbb{P}^{0}$ still resides. By shrinking the size of its ambiguity set, we expect to reduce the conservatism of DR-ARO. We start with the following assumption.
Assumption 5.1.
We are given finite $\varepsilon,\widehat{\varepsilon}>0$ such that $\mathrm{W}(\mathbb{P}^{0},\mathbb{P}_{N})\leq\varepsilon$ and $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}}_{\widehat{N}})\leq\widehat{\varepsilon}$.
We relax this assumption in Section 6. Given Assumption 5.1, we want to solve the revised problem:
@@ 5 Reducing conservatism of DR-ARO via intersection of Wasserstein balls
We first reformulate the intersected DR ARO problem (Inter-ARO) as a semi-infinite optimization problem with finite variables and then provide a complexity result.
Proposition 5.2.
Inter-ARO admits the following reformulation.
@@
$\begin{array}[]{cl}\underset{\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}}}{\inf}&\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\begin{aligned} \underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},l)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\leq s_{i}+\frac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\frac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda}\end{aligned}\\[14.22636pt] &\hfill\forall i\in[N],\ j\in[\widehat{N}],\ l\in\{-1,1\}\\ &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}.\end{array}$
@@
The proof is in Appendix B.3. Even though this problem recovers DR-ARO (hence admits tractable reformulations) when the radius $\widehat{\varepsilon}$ of the second ball satisfies $\widehat{\varepsilon}\rightarrow\infty$, Proposition 5.3 shows that it is NP-hard in the finite radius settings. We reformulate Inter-ARO as an adjustable robust optimization problem [5, 77], and borrow tools from this literature to obtain the following result.
Proposition 5.3.
@@
Inter-ARO is equivalent to an adjustable robust optimization problem with $\mathcal{O}(N\cdot\widehat{N})$ two-stage robust constraints, which is NP-hard even when $N=\widehat{N}=1$.
@@
The proof is in Appendix B.4. The adjustable robust optimization literature has developed a rich arsenal of relaxation techniques that can be leveraged for Inter-ARO. We adopt the ‘static relaxation technique’ [10] to restrict the feasible region of Inter-ARO and obtain a tractable approximation.
Theorem 5.4.
The following convex optimization problem is a feasible relaxation (safe approximation) of Inter-ARO:
@@
$\begin{array}[]{cl}\underset{\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}},\bm{z}_{ij}^{+},\bm{z}_{ij}^{-}}{\inf}&\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\left[\begin{aligned} &L^{\alpha}(l\cdot\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}_{ij}^{l\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))\leq s_{i}+\dfrac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\dfrac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda},\\[3.41432pt] &\lVert l\bm{\beta}-\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\lambda,\;\lVert\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\widehat{\lambda}\end{aligned}\right]\\[22.76219pt] &\hfill\forall i\in[N],\ j\in[\widehat{N}],\ l\in\{-1,1\}\\ &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+},\;\bm{z}^{l}_{ij}\in\mathbb{R}^{n},\end{array}$ (13)
where $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ is the univariate representation of $\ell_{\bm{\beta}}^{\alpha}$.
The proof is in Appendix B.5. 13 relaxes the NP-hard problem Inter-ARO so that it becomes efficiently solvable, and it enjoys similar tractable formulations to DR-ARO.
Remark 5.5.
13 admits an exponential cone reformulation, analogously to Appendix C.
@@
Recall that for $\widehat{\varepsilon}$ large enough so that $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})=\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$Inter-ARO reduces to DR-ARO. The following corollary (proof in Appendix B.6) shows that a similar desired property holds for the relaxed problem 13. That is, “not learning anything from auxiliary data” remains feasible: the static relaxation does not force learning from ^$\widehat{\mathbb{P}}_{\widehat{N}}$, it learns from auxiliary data only if the objective improves. Moreover, we show that as $\widehat{\varepsilon}\to\infty$13 converges to Inter-ARO.
Corollary 5.6.
@@
Feasibility of disregarding auxiliary data: Any feasible solution $(\bm{\beta},\lambda,\bm{s})$ of DR-ARO gives a feasible solution $(\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}},\bm{z}^{+}_{ij},\bm{z}^{-}_{ij})$ for 13 with $\widehat{\lambda}=0$, $\widehat{\bm{s}}=\mathbf{0}$, $\bm{z}^{+}_{ij}=\bm{z}^{-}_{ij}=\mathbf{0}$. Convergence to Inter-ARO: As $\widehat{\varepsilon}\rightarrow\infty$, the optimal value of 13 converges to the optimal value of Inter-ARO, with the same set of optimal $\bm{\beta}$ solutions.
Inter-ARO and Related Problems
@@
Recall that Inter-ARO can simply ignore the auxiliary data once $\widehat{\varepsilon}$ is set large enough, reducing this problem to DR-ARO. Moreover, notice that $\alpha=0$ reduces $\ell_{\bm{\beta}}^{\alpha}$ to $\ell_{\bm{\beta}}$, hence for $\alpha=0$ and $\widehat{\varepsilon}=\infty$ Inter-ARO recovers the Wasserstein LR model of [60]. We next relate Inter-ARO to the problems in the ARO literature that use auxiliary data $\{(\bm{\widehat{x}}^{j},\widehat{y}^{j})\}_{j\in[\widehat{N}]}$. The works in this literature [30, 76] propose solving the following for some $w>0$:
@@
$\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\ \frac{1}{N+w\widehat{N}}\big{[}\sum_{i\in[N]}\underset{\bm{z}^{i}\in\mathcal{B}_{p}(\alpha)}{\sup}\{\ell_{\bm{\beta}}(\bm{x}^{i}+\bm{z}^{i},y^{i})\}+w\sum_{j\in[\widehat{N}]}\underset{\bm{z}^{j}\in{\mathcal{B}_{p}(\alpha)}}{\sup}\{\ell_{\bm{\beta}}(\bm{\widehat{x}}^{j}+\bm{z}^{j},\widehat{y}^{j})\}\big{]},$ (14)
@@
Proposition 5.7.
Problem (14) is equivalent to:
$\begin{array}[]{cl}\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}&\mathbb{E}_{\mathbb{Q}_{\mathrm{mix}}}[\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)]\\ \end{array}$ (16)
@@
The proof is in Appendix B.7. Next, we give a condition on $\varepsilon$ and $\widehat{\varepsilon}$ to guarantee that the mixture distribution introduced in Proposition 5.7 lives in $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$, that is, the distribution $\mathbb{Q}_{\mathrm{mix}}$ will be feasible in the $\sup$ problem of Inter-ARO.
Proposition 5.8.
@@
For any $\lambda\in(0,1)$ and $\mathbb{Q}_{\mathrm{mix}}:=\lambda\cdot\mathbb{P}_{N}+(1-\lambda)\cdot\widehat{\mathbb{P}}_{\widehat{N}}$, whenever $\varepsilon+\widehat{\varepsilon}\geq\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$ and $\frac{\widehat{\varepsilon}}{\varepsilon}=\frac{\lambda}{1-\lambda}$ are satisfied, we have $\mathbb{Q}_{\mathrm{mix}}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$.

### v2

@@ 3 PRELIMINARIES
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\;\displaystyle\mathbb{E}_{{\mathbb{P}}_{N}}[\underset{\bm{z}:\lVert\bm{z}\rVert_{p}\leq\alpha}{\sup}\{\ell_{\bm{\beta}}(\bm{x}+\bm{z},y)\}].\end{array}$ (8)
@@
8 reduces to 4 when $\alpha=0$. Note that 8 is identical to feature robust training (Bertsimas et al.,, 2019) which is not motivated by adversarial attacks, but the presence of noisy observations in the training set (Ben-Tal et al.,, 2009; Gorissen et al.,, 2015).
DRO-ARO connection. A connection between ARO and DRO is noted in the literature (Staib and Jegelka, 2017, Proposition 3.1, Khim and Loh, 2018, Lemma 22, Pydi and Jog, 2021, Lemma 5.1, Regniez et al., 2022, Proposition 2.1, Frank and Niles-Weed, 2024, Lemma 3, and Bennouna et al., 2023, §3). Namely, problem 8 is equivalent to a DRO problem
@@
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\;\displaystyle\underset{\mathbb{Q}\in\mathfrak{B}^{\infty}_{\alpha}({\mathbb{P}}_{N})}{\sup}\;\mathbb{E}_{\mathbb{Q}}[\ell_{\bm{\beta}}(\bm{x},y)],\end{array}$ (10)
where the ambiguity set $\mathfrak{B}^{\infty}_{\alpha}({\mathbb{P}}_{N})$ is a type-$\infty$ Wasserstein ball (Givens and Shortt,, 1984) with radius $\alpha$. Hence, in non-adversarial settings, 8 provides robustness with respect to the type-$\infty$ Wasserstein distance. In the case of adversarial attacks, it suffers from robust overfitting as discussed earlier. To address this issue, one straightforward approach is to revisit (10) and replace $\alpha$ with some $\alpha^{\prime}>\alpha$. This approach, however, does not provide improvements for the out-of-sample performance since (i) the type-$\infty$ Wasserstein distance employed in problem (10) uses a metric on the feature space, ignoring labels; (ii) type-$\infty$ Wasserstein distances do not provide strong out-of-sample performances in ML (unlike, e.g., the type-$1$ Wasserstein distance) since the required radii to provide meaningful robustness guarantees are typically too large (Bennouna and Van Parys,, 2022, §1.2.2, and references therein). We thus study the type-$1$ Wasserstein counterpart of 8, which we initiate in the next section.
4 DISTRIBUTIONALLY AND ADVERSARIALLY ROBUST LR
Here we derive the Wasserstein DR counterpart of 8 that will set the ground for our main result in the next section. We impose the following assumption.
Assumption 1.
We are given a finite $\varepsilon>0$ value satisfying $\mathrm{W}(\mathbb{P}^{0},\mathbb{P}_{N})\leq\varepsilon$.
The assumption implies that we know an $\varepsilon>0$ value satisfying $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$. Typically, however, $\varepsilon$ is either estimated through cross-validation or finite sample statistics, with the assumption then regarded as holding with high confidence (see §6 for a review of related results we can borrow). The distributionally and adversarially robust LR problem is thus:
@@ 4 DISTRIBUTIONALLY AND ADVERSARIALLY ROBUST LR
By employing a simple duality trick for the inner $\sup$-problem, as commonly applied in robust optimization (Ben-Tal et al.,, 2009; Bertsimas and Den Hertog,, 2022), we can represent DR-ARO as a standard non-adversarial DRO problem with an updated loss function, which we name the adversarial loss.
Observation 1.
Problem DR-ARO is equivalent to
$\begin{array}[]{cl}\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\underset{\mathbb{Q}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})}{\sup}\mathbb{E}_{\mathbb{Q}}[\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)],\end{array}$
where the adversarial loss $\ell^{\alpha}_{\bm{\beta}}$ is defined as
$\ell^{\alpha}_{\bm{\beta}}(\bm{x},y):=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}})),$
for $p^{\star}$ satisfying $11/p+1/p^{\star}=$ 1. The univariate representation $L^{\alpha}(z):=\log(1+\exp(-z+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))$ of $\ell^{\alpha}_{\bm{\beta}}$ is convex and has a Lipschitz modulus of $1$.
As a corollary of Observation 1, we can directly employ the techniques proposed by Shafieezadeh-Abadeh et al., (2019) to dualize the inner $\sup$-problem of DR-ARO and obtain a tractable reformulation.
Corollary 1.
@@
$\begin{array}[]{cll}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\bm{s}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\dfrac{1}{N}\sum_{i=1}^{N}s_{i}\\[11.99998pt] \mathrm{s.t.}&\displaystyle\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},y^{i})\leq s_{i}&\forall i\in[N]\\[1.99997pt] &\displaystyle\ell_{\bm{\beta}}^{\alpha}(\bm{x}^{i},-y^{i})-\lambda\kappa\leq s_{i}&\forall i\in[N]\\[1.99997pt] &\displaystyle\lVert\bm{\beta}\rVert_{q^{\star}}\leq\lambda\\[1.99997pt] &\displaystyle\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\end{array}$
for $q^{\star}$ satisfying $11/q+1/q^{\star}=$ 1.
The constraints of this problem are exponential cone representable (derivation is in the appendices) and for $q\in\{1,2,\infty\}$, the yielding problem can be solved with the exponential cone solver of MOSEK (MOSEK ApS,, 2023) in polynomial time (Nesterov,, 2018).
5 MAIN RESULT
In §4 we discussed the traditional DRO setting where we have access to an empirical distribution $\mathbb{P}_{N}$ constructed from $N$ i.i.d. samples of the true data-generating distribution $\mathbb{P}^{0}$, and we are given (or we estimate) some $\varepsilon$ so that $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$. Recently in DRO literature, it became a key focus to study the case where we have access to an additional auxiliary empirical distribution ^$\widehat{\mathbb{P}}_{\widehat{N}}$ constructed from $\widehat{N}$ i.i.d. samples $\widehat{\bm{\xi}}^{j}=(\widehat{\bm{x}}^{j},\widehat{y}^{j}),\ j\in[\widehat{N}]$, of some other distribution $\widehat{\mathbb{P}}$; given the increasing availability of useful auxiliary data in the ARO domain, we explore this direction here. We start with the following assumption.
Assumption 2.
We are given finite $\varepsilon,\widehat{\varepsilon}>0$ values satisfying $\mathrm{W}(\mathbb{P}^{0},\mathbb{P}_{N})\leq\varepsilon$ and $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}}_{\widehat{N}})\leq\widehat{\varepsilon}$.
The assumption implies that we know $\varepsilon,\widehat{\varepsilon}>0$ values satisfying $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$. In practice, this assumption is ensured to hold with high confidence by estimating the $\varepsilon$ and $\widehat{\varepsilon}$ values; methods across various domains which we can adopt are reviewed in §6. Under Assumption 2, we want to optimize the adversarial loss over the intersection $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$:
@@ 5 MAIN RESULT
This formulation is expected to outperform DR-ARO as the ambiguity set is smaller while still including $\mathbb{P}^{0}$. However, problem Inter-ARO is challenging to solve even in the absence of adversarial attacks ($\alpha=0$) as we reviewed in §2. To address this challenge, we first reformulate Inter-ARO as a semi-infinite optimization problem with finitely many variables.
Proposition 1.
Inter-ARO is equivalent to:
@@
$\begin{array}[]{cl}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\widehat{\lambda}\\ \bm{s},\widehat{\bm{s}}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\left[\begin{aligned} \underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\{\ell^{\alpha}_{\bm{\beta}}(\bm{x},l)-\lambda\lVert\bm{x}^{i}-\bm{x}\rVert_{q}-\widehat{\lambda}\lVert\widehat{\bm{x}}^{j}-\bm{x}\rVert_{q}\}\\ \leq s_{i}+\frac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\frac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda}\end{aligned}\right]\\[14.22636pt] &\hfill\forall i\in[N],\ j\in[\widehat{N}],\ l\in\{-1,1\}\\[2.84526pt] &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}.\end{array}$
@@
Even though this problem recovers the tractable problem DR-ARO as $\widehat{\varepsilon}\rightarrow\infty$, it is NP-hard in the finite radius settings. We reformulate Inter-ARO as an adjustable robust optimization problem (Ben-Tal et al.,, 2004; Yanıkoğlu et al.,, 2019), and borrow tools from this literature to obtain the following result.
Proposition 2.
@@
Inter-ARO is equivalent to an adjustable RO problem with $\mathcal{O}(N\cdot\widehat{N})$ two-stage robust constraints, which is NP-hard even when $N=\widehat{N}=1$.
@@
The adjustable RO literature has developed a rich arsenal of relaxations that can be leveraged for Inter-ARO. We adopt the ‘static relaxation technique’ (Bertsimas et al.,, 2015) to restrict the feasible region of Inter-ARO and obtain a tractable approximation.
Theorem 1 (main).
The following convex optimization problem is a feasible relaxation of Inter-ARO:
@@
$\begin{array}[]{cl}\underset{\begin{subarray}{c}\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}}\\ \bm{z}_{ij}^{+},\bm{z}_{ij}^{-}\end{subarray}}{\inf}&\displaystyle\varepsilon\lambda+\widehat{\varepsilon}\widehat{\lambda}+\frac{1}{N}\sum_{i=1}^{N}s_{i}+\frac{1}{\widehat{N}}\sum_{j=1}^{\widehat{N}}\widehat{s}_{j}\\ \mathrm{s.t.}&\left[\begin{aligned} &L^{\alpha}(l\cdot\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}_{ij}^{l\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))\\ &\ \leq s_{i}+\dfrac{\kappa(1-ly^{i})}{2}\lambda+\widehat{s}_{j}+\dfrac{\kappa(1-l\widehat{y}^{j})}{2}\widehat{\lambda},\\[3.41432pt] &\lVert l\bm{\beta}-\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\lambda,\;\lVert\bm{z}^{l}_{ij}\rVert_{q^{\star}}\leq\widehat{\lambda}\end{aligned}\right]\\[22.76219pt] &\hfill\forall i\in[N],\ j\in[\widehat{N}],\ l\in\{-1,1\}\\ &\bm{\beta}\in\mathbb{R}^{n},\;\lambda\geq 0,\;\widehat{\lambda}\geq 0,\;\bm{s}\in\mathbb{R}^{N}_{+},\;\widehat{\bm{s}}\in\mathbb{R}^{\widehat{N}}_{+}\\ &\bm{z}^{l}_{ij}\in\mathbb{R}^{n},\ (i,j,l)\in[N]\times[\widehat{N}]\times\{-1,1\}.\end{array}$ (16)
Similarly to DR-ARO, the constraints of 16 are exponential cone representable (cf. appendices).
@@
Recall that for $\widehat{\varepsilon}$ large enough, Inter-ARO reduces to DR-ARO. The following corollary shows that, despite 16 being a relaxation of Inter-ARO, a similar property holds. That is, “not learning anything from auxiliary data” remains feasible: the static relaxation does not force learning from ^$\widehat{\mathbb{P}}_{\widehat{N}}$, and it learns from auxiliary data only if the objective improves.
Corollary 2.
@@
Feasibility of ignoring auxiliary data: Any feasible solution $(\bm{\beta},\lambda,\bm{s})$ of DR-ARO can be used to recover a feasible solution $(\bm{\beta},\lambda,\widehat{\lambda},\bm{s},\widehat{\bm{s}},\bm{z}^{+}_{ij},\bm{z}^{-}_{ij})$ for 16 with $\widehat{\lambda}=0$, $\widehat{\bm{s}}=\mathbf{0}$, and $\bm{z}^{+}_{ij}=\bm{z}^{-}_{ij}=\mathbf{0}$.
Convergence to Inter-ARO: The optimal value of 16 converges to the optimal value of Inter-ARO, with the same set of $\bm{\beta}$ solutions, as $\widehat{\varepsilon}\rightarrow\infty$.
@@
We close the section by discussing how Inter-ARO can recover some problems in the DRO and ARO literature. Firstly, recall that Inter-ARO can ignore the auxiliary data once $\widehat{\varepsilon}$ is set large enough, reducing this problem to DR-ARO. Moreover, notice that $\alpha=0$ reduces $\ell_{\bm{\beta}}^{\alpha}$ to $\ell_{\bm{\beta}}$, hence for $\alpha=0$ and $\widehat{\varepsilon}=\infty$ Inter-ARO recovers the Wasserstein LR model of Shafieezadeh-Abadeh et al., (2015). We next relate Inter-ARO to the problems in the ARO literature that use auxiliary data. The works in this literature (Gowal et al.,, 2021; Xing et al., 2022b, ) solve the following for some $w>0$:
@@
$\begin{array}[]{r}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}\ \dfrac{1}{N+w\widehat{N}}\big{[}\sum_{i\in[N]}\underset{\bm{z}^{i}\in\mathcal{B}_{p}(\alpha)}{\sup}\{\ell_{\bm{\beta}}(\bm{x}^{i}+\bm{z}^{i},y^{i})\}+\\ \displaystyle w\sum_{j\in[\widehat{N}]}\underset{\bm{z}^{j}\in{\mathcal{B}_{p}(\alpha)}}{\sup}\{\ell_{\bm{\beta}}(\bm{\widehat{x}}^{j}+\bm{z}^{j},\widehat{y}^{j})\}\big{]},\end{array}$ (19)
@@
Observation 2.
Problem (19) is equivalent to
$\begin{array}[]{cl}\displaystyle\underset{\bm{\beta}\in\mathbb{R}^{n}}{\inf}&\displaystyle\mathbb{E}_{\mathbb{Q}_{\mathrm{mix}}}[\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)]\end{array}$ (21)
@@
We give a condition on $\varepsilon$ and $\widehat{\varepsilon}$ to guarantee that the mixture distribution introduced in Proposition 2 lives in $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$, that is, the distribution $\mathbb{Q}_{\mathrm{mix}}$ will be feasible in the $\sup$ problem of Inter-ARO.
Proposition 3.
@@
For any $\lambda\in(0,1)$ and $\mathbb{Q}_{\mathrm{mix}}=\lambda\cdot\mathbb{P}_{N}+(1-\lambda)\cdot\widehat{\mathbb{P}}_{\widehat{N}}$, we have $\mathbb{Q}_{\mathrm{mix}}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ whenever $\varepsilon+\widehat{\varepsilon}\geq\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$ and $\frac{\widehat{\varepsilon}}{\varepsilon}=\frac{\lambda}{1-\lambda}$.